The noise level in a classroom in absence of the teacher is 50 dB when 50 student s are present. Assuming that on the average each student outputs same sound energy per second, what will be the noise level if the number of students is increased to 100?

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
Sol. Let the intensity of each student be I. According to the question β base A = 10log base 10 50 l/l base 0 ; β base B = 10 log (100 l/l base 0) = 10 l log(100 l/50 l) = 10 log base 10 2 = 3 So, β base A = 50 + 3 = 53 dB.